OpenStudy (anonymous):

find k so that the line through 4,k and 2,-1 is perpendicular to y=2/5x-3

OpenStudy (anonymous):

the slope of the line is -5. rise/run=-5 (k+1)/(4-2)=-5 solve for k

OpenStudy (anonymous):

i mean -5/2 is the slope

OpenStudy (anonymous):

do you know how to finish the problem?

OpenStudy (anonymous):

no.

OpenStudy (anonymous):

\[\frac{-5}{2}=\frac{k+1}{4-2}\]now "cross multiply" this clears the 2 from the denominator\[-5=k+1\]subtract 1 to find k\[-6=k\]

OpenStudy (anonymous):

no you did it wrong.....shouldn't it be y2-y1 instead of y1-y2

OpenStudy (anonymous):

so instead of k+1 at the top it should be -1-k

OpenStudy (anonymous):

the slope of the given line is 2/5 the slope of the required line is -5/2 so the line must be in the form ,y=(-5/2)x+C,where C is some constant now apply the given points, (2,-1) we get, -1=(-5/2) (2)+C =>C=4 so the required line equation is y=(-5/2)x+4 now apply the other point ,(4,k)=>k=(-5/2)4+4 =>k=-10+4 =>k=-6

OpenStudy (anonymous):

what if the points were written (2,-1) and (4,k)? would k then be y2?

OpenStudy (anonymous):

yeah, but it's written as 4,k and 2,-1...therfore -1 is y2 and u start with it.

OpenStudy (anonymous):

not so, what matters in the approach that colton started, is that you remain consistent with respect to the position of coordinates when you set-up the slope formula. Simply because one point is written to the left of another in the problem does not make it point1 or point 2, that is you choice. I chose the set that gave the fewest number of negative signs

OpenStudy (anonymous):

but shouldnt we then get the same answer

OpenStudy (anonymous):

is went right to left and i got -12

OpenStudy (anonymous):

i**

OpenStudy (anonymous):

That said, if you set it up and did it according to \[\frac{-5}{2}=\frac{-1-y}{2-4}\]you would get the same result.

OpenStudy (anonymous):

i did that, then i got -10=-k-2

OpenStudy (anonymous):

-10+20 i mean

OpenStudy (anonymous):

you forgot to divide by 2 of the LHS; the LHS should be -5 not -10, then it works out

OpenStudy (anonymous):

???

OpenStudy (anonymous):

lhs?

OpenStudy (anonymous):

LHS=left hand side

OpenStudy (anonymous):

dont u cross multiply and dont have to divide?

OpenStudy (anonymous):

lets see, if we cross multiply we get\[-2*(-5)=2(-1-k)\]\[10=-2-2k\]\[12=-2k\]\[-6=k\]

OpenStudy (anonymous):

you forgot +20 in the first step and completely ignored the -4

OpenStudy (anonymous):

it's 2-4

OpenStudy (anonymous):

also go see dashni's solution; you might like that method better

OpenStudy (anonymous):

2-4 in the right side denominator is -2; then you see the negative mutliplying the -5 on the right side after the cross mult.

OpenStudy (anonymous):

i dont get why u just ignmored the -4 though

OpenStudy (anonymous):

i didn't ignore the -4; i simplified 2-4=-2!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!

OpenStudy (anonymous):

ohhhhhhhhhhhhhhhhhhhhhhhh

OpenStudy (anonymous):

WOW THAT MAKES SENSE NOW

OpenStudy (anonymous):

there were a few responses between the start and restart of that way; I should have started from the beginning; btw, go see dashni's solution to the problem-- you might find it more to your liking

OpenStudy (anonymous):

no i didnt get tht...urs was better.

OpenStudy (anonymous):

actually it was colton's to start

OpenStudy (anonymous):

in other news....have u read the hunger games?

OpenStudy (anonymous):

I might check it out, thanks!